OpenStudy (anonymous):

Write an equation of the line that passes through the given point and is parallel to the given line. (9,4)y-x=3

OpenStudy (saifoo.khan):

y = x +3 Slope is 1.

OpenStudy (saifoo.khan):

Now, the line which will be parallel will have the same slope as well.

OpenStudy (saifoo.khan):

So, \[x_2 - x_1 = m(y_2 - y_1)\]

OpenStudy (anonymous):

k i see

OpenStudy (saifoo.khan):

x - 9 = 1(y - 4) Solve. :)